OpenStudy (anonymous):

solve and find the value of x . 3x^2+78=4x^2+7x

OpenStudy (jiteshmeghwal9):

4x^2-3x^2+7x-78=0 x^2+7x-78=0 x^2+13x-6x-78 x(x+13)-6(x+13)=0 (x-6)(x+13)=0 x=6 & x=-13

OpenStudy (ajprincess):

\(3x^2+78=4x^2+7x\) \(4x^2+7x-3x^2-78=0\) \(x^2+7x-78=0\) use quadratic formula to find x \[x=\frac{-b\pm\sqrt{b^2-4ac} }{ 2a }\] Here a=1 b=7 c=-78